# Palu: Compressing KV-Cache with Low-Rank Projection

### v1

@@ 3.2.3 Group-Head Low-Rank Decomposition
The FLOPs for reconstructing the keys and values for each head in G-LRD is $r_{g}\cdot d_{h}\cdot n_{g}$, where $n_{g}=\frac{n}{s}$ is the number of groups. Comparing the cost to J-LRD and assuming the same total rank size ($r_{g}\cdot n_{g}=r_{\text{joint}}$), G-LRD reduces the reconstruction cost by $n_{g}$. Similarly, G-LRD also reduces the fused matrix size by $n_{g}$. To sum up, G-LRD offers a middle ground between computation overhead and approximation accuracy. We illustrate M-LRD, J-LRD and G-LRD in Fig.3.
For further discussion on decomposed weight matrix size, please refer to Appendix B, where more detailed information is provided.
@@ 3.3 Automatic Rank Allocation
To allocate an ideal rank size to the decomposition target, it is crucial to accurately estimate the importance of the target matrix (e.g., grouped weights). In Palu, we identify Fisher information (Ly et al., 2017; Liu et al., 2021) as an accurate approximator since it can quantify the amount of information for each parameter. We then employ the sum of Fisher information to estimate the importance of the weight matrix of each linear layer.
Assuming that the compression sensitivity is proportional to Fisher information, we determine the rank for each weight matrix by computing the ratio of its Fisher information to the total Fisher information across all decomposition targets. We use this ratio to allocate the compression rate (i.e., rank level $r$), ensuring that more important layers retain higher rank levels.
For a detailed ablation study on our Automatic Rank Allocation method, please refer to Appendix H.3, where we analyze the impact of our strategy on model performance.
3.4 Low-Rank-Aware Quantization
Figure 4: Activation distribution of the low-rank compressed key caches at the $4^{th}$ Llama-2 attention layer.
We integrate quantization into Palu to further compress the KV-Cache. However, we observe that low-rank compressed latent representations have severe outliers, which limit quantization applicability in Palu. Unlike natural outliers described in previous KV-Cache quantization literature (Liu et al., 2024; Hooper et al., 2024), these outliers are induced by SVD-based low-rank factorization.
@@ 3.4 Low-Rank-Aware Quantization
Inspired by recent LLM quantization literature (Ashkboos et al., 2024b; Tseng et al., 2024), we apply the Walsh-Hadamard transform (WHT, Fino and Algazi) to eliminate outliers (Fig. 4 (b)), enabling a high quantization accuracy. However, this transformation introduces an extra matrix multiplication with associated runtime overhead. Unlike earlier methods (Ashkboos et al., 2024b) that must apply online WHT when quantizing KV-Cache, we optimize this process by integrating the Hadamard matrix into low-rank decomposed weights with no additional overhead, as described by
@@
where $\mathbf{R}$ is the Hadamard matrix. This optimization allows Palu can integrate the proposed low-rank compression technique with low-bit quantization. Our experiments show that, on top of the low-rank compression, our quantization method only negligibly increases perplexity, even at extreme levels such as 3-bit or 2-bit (see Sec.4.2).
3.5 Efficient Inference Implementation
Support for RoPE-based Attention.
To support RoPE-based attention, we fuse the reconstruction of original keys, applying positional embedding (i.e., RoPE (Su et al., 2024)) and the final matrix-vector multiplication of query and reconstructed keys by implementing customized GPU kernels in Triton Tillet et al. (2019). We leverage the tiling techniques akin to FlashAttention (Dao et al., 2022) to minimize the memory reads/writes of GPU high bandwidth memory.
Table 1: Perplexity and zero-shot accuracy of Palu, with different decomposition strategies at 50% compression rate.
@@ 3.5 Efficient Inference Implementation
Llama2-13B Baseline 4.88 6.70 45.20 79.39 79.11 79.42 49.06 72.38 67.43
J-LRD 4.97 6.92 46.40 79.48 78.62 79.29 49.91 70.56 67.38
G-LRD 5.31 7.76 45.60 77.29 77.42 76.05 45.99 72.45 65.80
M-LRD 5.65 8.34 43.20 74.34 77.53 75.76 45.39 68.98 64.20
4 Experiments
Table 2: Accuracy of Palu with 30% and 50% compression rate on 8 tasks in LongBench.
Model Method SAMSum LCC M-News Qasper TREC QmSum R-Bench TriQA Avg.
Vicuna-7B Baseline 40.73 49.20 27.98 25.57 69.00 22.54 43.50 86.42 45.62
Palu-30% 40.42 49.00 26.96 18.54 65.50 21.78 45.77 85.22 44.15
Palu-50% 38.36 34.59 24.78 19.77 63.50 21.44 35.42 79.74 39.70
Mistral-7B Baseline 41.71 66.59 20.26 8.13 67.00 19.99 59.10 89.90 46.59
Palu-30% 43.23 64.23 23.50 10.21 72.50 21.16 56.80 89.82 47.68
Palu-50% 36.43 50.22 19.23 7.54 67.50 19.56 44.76 85.37 41.33
Models and Tasks.
We evaluate Palu on three LLM families, Llama2(Touvron et al., 2023), Mistral(Jiang et al., 2023) and Vicuna(Chiang et al., 2023). For accuracy evaluation, we measure perplexity on the WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020) datasets and use LM-Evaluation-Harness (Gao et al., 2023) to measure zero-shot accuracy on six common sense tasks. We also evaluate long context accuracy on 8 tasks in LongBench (Bai et al., 2023). See Appendix I for further details on the dataset and settings.
Compression Settings.
We implemented Palu using the Huggingface library Wolf et al. (2020). Decomposition of the key and value projection layers was performed using the truncation-aware SVD method proposed by SVD-LLM (Wang et al., 2024). Unless otherwise specified, results for Palu are reported using G-LRD with a group size (gs) of 4 with equal rank size for each group. To calculate Fisher information in rank searching, we used 2048 samples from Wikitext-2, each with a sequence length of 1024. For quantization, we aligned our settings with Atom (Zhao et al., 2023), and use a simple per-token, asymmetric quantization to quantize the low-rank latent representation of both key and value states.
4.1 Main Results
@@
We first present perplexity evaluation results on Wikitext-2 and C4 in Table 1. For the Llama2-7B model, we observe that the M-LRD method fails to maintain low perplexity at a 50% compression rate. In contrast, the J-LRD method performs significantly better given the same compression rate, achieving perplexity values of 5.62 on Wikitext-2. The G-LRD method, with a group size of 4, achieves an optimal balance between perplexity and computational cost, yielding perplexity values of 6.01 on Wikitext-2 at a 50% compression rate. The same trend is consistent across the Llama2-13B model.
@@
We further evaluate Palu on six common-sense zero-shot tasks, as shown in Table 1. Similar to the perplexity evaluation, the J-LRD method demonstrates the best performance on Llama2-7B, with only a 0.23% average accuracy degradation, while the M-LRD method results in the lowest average performance, with a 7.26% drop in accuracy compared to the baseline. In comparison, G-LRD experiences only a 2.4% average accuracy decline.
LongBench Results.
To assess whether Palu can effectively compress the KV-Cache for long-context scenarios, we present the performance of Palu across eight tasks from LongBench. We select Mistral-7B and Vicuna-7B (Chiang et al., 2023) as our evaluation targets, which support sequences up to 8k and 16k, respectively. Table 2 reports experiment results. We find that at a 50% compression level, it is relatively difficult to fully preserve accuracy. However, when compressing KV-Cache for 30%, Palu only has a 1.47% averaged accuracy degradation for Vicuna-7B, and even achieves a 1.09% enhancement for Mistral-7B. These results evidence of Palu’s ability to shrink the KV-Cache while maintaining performance.
For detailed information on the integration of Palu with quantization and the performance measurements on LongBench, please refer to Appendix E.
4.2 Integrating Palu with Quantization
Table 3 shows the results of integrating quantization with Palu. With 3-bit quantization, Palu exhibits only a negligible perplexity increase of 0.08 at a 30% compression rate and a minimal 0.23 at a 50% compression rate, both compared to Palu at 16-bit. This indicates a minimal trade-off for substantial compression of integrating quantization. Furthermore, at 2-bit quantization, Palu significantly outperforms KVQuant, which is a SOTA quantization-based method. Palu demonstrates a compelling reduction in perplexity by 1.19 and 0.54 points, while consuming 30% and 50% less memory, respectively. These results underscore the effectiveness of Palu, establishing it as a superior KV-Cache compression strategy.
4.3 Latency Evaluation
We now provide the latency evaluation for Palu. In our experiment, latency is measured on an A6000 GPU, and the averaged results over 100 runs are reported. For baseline, we use the default implementation from HuggingFace in PyTorch. We do not apply FlashAttention (Dao et al., 2022) for an apple-to-apple comparison, as our current implementation does not fuse the Softmax operation. We leave the integration with FlashAttention as the future work.
GEMV with Online Reconstruction.
We evaluate the performance of our fused reconstruction kernel (Sec. 3.5) for computing RoPE-based attention score. Specifically, we measure the latency from pre-RoPE query vector to the post-GEMV attention score (refer to Fig. 2 for details). Here, we set the compression rate for our kernel as 75%. It’s because we observe the key states often achieve such compression rate when setting a 50% overall compression rate (see Appendix H.3 for detailed distribution).
We compare Palu’s runtime against the baseline, which is simply a GEMV, and report the speedup in Fig. 5 (a). For group size 32 (J-LRD), although it’s the most accurate decomposition, the high reconstruction leads to a significant 7.09$\times$ slowdown at sequence length 64K . For group size 1 (M-LRD), our fused kernel achieves up to a 2.48$\times$ speedup at a sequence length 16k, showcasing a high speedup when more accuracy loss is acceptable. For group size 2 and 4 (G-LRD), our fused kernel demonstrates 1.75$\times$ and 1.16$\times$ speedup at sequence length 16K. These results highlight the importance of exploring trade-off between decomposition granularity, reconstruction cost, and accuracy. Furthermore, we observe our kernel speedup decreases for sequence lengths over 16k, due to the increasing reconstruction cost shifts the attention operation from memory- to compute-bounded.
Table 3: Perplexity and KV-Cache size comparison of Palu and other quantization-based methods for Llama2-7B on WikiText-2. For perplexity, sequence length is set to 4096, aligning with baselines. KV-Cache size is demonstrated for a sequence length of 128K.
@@ 4.3 Latency Evaluation
End-to-end attention module.
We measure the end-to-end latency of the entire RoPE-based attention module, with the integration of our fused reconstruction kernel for key states (Sec. 3.5) and offline matrix fusion for value states (Eq. 5). Here, we set the compression rate for key and value as 75% and 25%, for an overall 50% compression rate. The detailed forward pass operations are presented in Fig. 2. We measure the decoding runtime for sequence length from 4K to 64K, and report the runtime of Palu and baseline in Fig. 5 (b). As Fig. 5 (b), Palu with smaller group sizes consistently delivers more than 1.43$\times$ speedup across all sequence lengths. For group size 1 (M-LRD), Palu achieves a 1.98$\times$, close to optimal (for 50% compression rate), speedup for sequence length 16K. For group size 4 (G-LRD), which is the sweet spot for maintaining strong accuracy and limiting reconstruction cost, Palu achieves up to 1.61$\times$ end-to-end speedup at sequence length 64K.
We observe increasing speedup as sequence length increases from 4k to 64k due to continuously reducing latency from compressing value states. However, speedup is expected to converge or slightly decrease with further length increases due to higher reconstruction costs and the compute-bound nature of reconstruction kernels. We expect the speedup of Palu will be larger for non-RoPE-based attention, as the reconstruction overhead can be entirely eliminated by applying Eq. 6.
For the end-to-end model latency, please refer to Appendix D, where more detailed information and data are provided.
Figure 5: Normalized speedup for (a) reconstruction kernel at 75% compression and (b) RoPE-based attention module with an overall 50% compression under different group sizes.
5 Related Work
SVD for LLM Compression.
There are several related works applying SVD for LLM compression. An early work (Noach and Goldberg, 2020) simply compresses the weight matrices using standard SVD, resulting in high compression errors. To address this, FWSVD (Hsu et al., 2022) incorporates Fisher information to prioritize parameters. Additionally, ASVD (Yuan et al., 2023) taken the activation outlier phenomenon in LLM into consideration, and propose and scaling techniques to mitigate the compression error. SVD-LLM (Wang et al., 2024) further advances this work by understanding the compression loss from each singular value, improving compression accuracy. Unlike previous related SVD works that focus on compressing model weights, Palu focus on using SVD to reduce KV-Cache size.
KV-Cache Quantization.
Quantization appears to be an effective compression technique to compress KV-Cache. Atom (Zhao et al., 2023) uses a simple per-token quantization scheme to compress KV-Cache. WKVQuant (Yue et al., 2024) proposes a two-level quantization scheme to improve quantization accuracy. KIVI (Liu et al., 2024) utilizes different quantization schemes, per-channel and per-token, for key and value states respectively. KVQuant (Hooper et al., 2024) follow a similar setting but adopts non-uniform quantization and sparse matrix for preserving outliers. In Palu, we eliminate outliers with WHT to enhance quantization friendliness and, therefore, only need a simple per-token quantization and achieve a remarkable accuracy.
6 Conclusion
We introduce Palu, a novel KV-Cache compression framework that decomposes key and value linear projection weight matrices and caches the compressed latent representations. We propose various optimizations, including group-head low-rank decomposition, rank search algorithm, low-rank-aware quantization, matrix fusion, and kernel fusion. With these optimizations, Palu can maintain accuracy while achieving significant memory reduction and faster inference. Experiments show that Palu can accelerate the end-to-end attention module by up to 1.61$\times$ with 50% KV-Cache compression; when combined with 2-bit quantization, Palu achieves a remarkable 91.25% compression rate, both maintaining strong accuracy.
Limitations and Future Work
We introduce a novel KV-Cache compression framework based on low-rank projection. However, our experiments cover model variants with up to only 13B parameters due to limited computational resources. Future work needs to establish how Palu performs for models with larger sizes (e.g., 70B or beyond). Also, we leave for future work the evaluation of end-to-end speedup when combining low-rank compression and quantization. It would also be interesting to investigate how to combine Palu with other efficient LLM methods, such as token eviction or weight quantization, to achieve the highest possible acceleration.
Acknowledgement
We would like to thank Sandy Kaplan for her valuable feedback and support in refining the writing of this paper. Her insights and suggestions significantly contributed to the clarity and quality of the final manuscript.
Appendix
Appendix A Quantization Basics
[…]
Appendix B Discussion regarding memory usage
This issue primarily arises in J-LRD decomposition schemes, where the projections of all heads are decomposed jointly. In contrast, our M-LRD decomposition schemes and optimized G-LRD schemes involve non-square target matrices. For example, in the G-LRD scheme with a group size of 4, the target matrix is formed by concatenating the original projection matrices of each attention head in the group. In the Llama-2-7b model, with an embedding dimension of 4096 and head dimensions of 128, each projection matrix is 4096x128, resulting in a concatenated matrix of size 4096x512. In this case, the dimensions should be considered as $m=8n$. Applying the referenced equation $\frac{mr+nr}{mn}$ with $r=0.7n$, we find that $\frac{mr+nr}{mn}=0.7875$, indicating no additional storage cost and, in fact, achieving an additional 21.25% memory savings.
Appendix C Overall Memory Reduction
Figure 6: Total memory usage for various sequence lengths in Llama2-7B. low-rank compression is at 50% for Palu.
Appendix D End-to-End Latency Measurements
We provide the latency evaluation of integrating our optimized attention modules into the Llama-2 models in Table 4. Measurement times are expressed in milliseconds (ms). Our measurements indicated that Palu achieved significant end-to-end speedup during the decoding phase, especially with longer sequence lengths. Specifically, for a sequence length of 32K, the speedups were 1.40x, 1.55x, and 1.71x for gs=4, gs=2, and gs=1, respectively. The results demonstrate the effectiveness of our methods and the potential to explore the redundancy in the hidden dimension.
Table 4: End-to-end decoding latency of Llama-2-7B with Palu measured on A6000.
Method Group Sequence Length
Size 4k 8k 16k 32k
Baseline - 36.07 44.03 69.28 112.75
J-LRD 32 86.39 103.41 140.85 OOM
G-LRD 4 32.12 38.89 52.69 80.81
G-LRD 2 28.82 34.94 47.45 72.89
M-LRD 1 27.00 32.46 43.56 66.13
Appendix E Extended Experiments on LongBench and comparisons to quantization baselines
We conducted extensive experiments using the Mistral-Instruct-7B-v0.2 (Jiang et al., 2023) and LongChat-7b-v1.5-32k (Li et al., 2023) models, both with extended context lengths of 32k tokens. We included 16 available English datasets from LongBench to ensure a comprehensive evaluation. We report the average score for each type of task separately, as well as the overall average across all 16 tasks. This evaluation aimed to broadly assess the effectiveness of Palu in long-context scenarios. For reference, we also include comparisons with KIVI.
Our findings indicate that at a 30% compression level, Palu achieves only a minor average accuracy drop (<1%) compared to the baseline for both models. These impressive results are attributed to our effective use of low-rank projection methods to explore redundancy in hidden dimensions. Furthermore, by fusing the Hadamard matrix offline to improve quantization friendliness, we were able to quantize the low-rank projected KV-Cache down to 3 bits. This quantization process resulted in less than a 1% accuracy degradation, achieved with a straightforward per-token quantization approach. When compared to KIVI, our framework demonstrates the ability to harness an additional 30% of redundancy. Notably, Palu does not require the complex grouped quantization and mixed-precision techniques employed by KIVI, highlighting the quantization friendliness that Palu offers.

### v2

@@ 3.2.3 Group-Head Low-Rank Decomposition
The FLOPs for reconstructing the keys and values for each head in G-LRD is $r_{g}\cdot d_{h}\cdot n_{g}$, where $n_{g}=\frac{n}{s}$ is the number of groups. Comparing the cost to J-LRD and assuming the same total rank size ($r_{g}\cdot n_{g}=r_{\text{joint}}$), G-LRD reduces the reconstruction cost by $n_{g}$. Similarly, G-LRD also reduces the fused matrix size by $n_{g}$. To sum up, G-LRD offers a middle ground between computation overhead and approximation accuracy. We illustrate M-LRD, J-LRD and G-LRD in Fig. 3. Please refer to Appendix C for further discussions on the costs of different decomposition granularities.
@@ 3.3 Automatic Rank Allocation
To allocate an ideal rank size to the decomposition target, it is crucial to accurately estimate the importance of the target matrix (e.g., grouped weights). In Palu, we identify Fisher information (Ly et al., 2017; Liu et al., 2021) as an accurate approximator since it can quantify the amount of information for each parameter. We then employ the sum of Fisher information to estimate the importance of the weight matrix of each linear layer (Abdelfattah et al., 2021).
Assuming that the compression sensitivity is proportional to Fisher information, we determine the rank for each weight matrix by computing the ratio of its Fisher information to the total Fisher information across all decomposition targets. We use this ratio to allocate the compression rate (i.e., rank level $r$), ensuring that more important layers retain higher rank levels. For a detailed ablation study on our Automatic Rank Allocation, please refer to Appendix G.3.
3.4 Quantization Compatibility
Figure 4: Activation distribution of the low-rank key caches at the $4^{th}$ Llama-2 attention layer.
We integrate quantization into Palu to compress the KV-Cache further. We observe that low-rank compressed latent representations have severe outliers, which limit quantization applicability in Palu. Unlike natural outliers described in previous KV-Cache quantization literature (Liu et al., 2024b; Hooper et al., 2024), these outliers are induced by SVD-based low-rank factorization.
@@ 3.4 Quantization Compatibility
Inspired by recent LLM quantization literature (Ashkboos et al., 2024b; Tseng et al., 2024), we apply the Walsh-Hadamard transform (WHT, Fino & Algazi) to eliminate outliers (Fig. 4 (b)), enabling a high quantization accuracy. However, this transformation introduces an extra matrix multiplication with associated runtime overhead. Unlike earlier methods (Ashkboos et al., 2024b) that must apply online WHT when quantizing KV-Cache, we optimize this process by integrating the Hadamard matrix into low-rank decomposed weights with no additional compute overhead, as described by
@@
where $\mathbf{R}$ is the Hadamard matrix. This optimization allows Palu to integrate the proposed low-rank compression technique with low-bit quantization. Our experiments show that, on top of the low-rank compression, our quantization method only negligibly increases perplexity, even at extreme levels such as 3-bit or 2-bit with a simple per-token quantization scheme (see Sec. 4.3).
4 Experiments
4.1 Experiments Setup
Models and Tasks.
We evaluate Palu on four LLM families, Llama-2 (Touvron et al., 2023), Llama-3 (Dubey et al., 2024), Mistral (Jiang et al., 2023) and LongChat (Li et al., 2023). For accuracy evaluation, we measure perplexity on the WikiText-2 (Merity et al., 2016) and C4 (Raffel et al., 2020) datasets and use LM-Evaluation-Harness (Gao et al., 2023) to measure zero-shot accuracy on six common sense tasks. We also evaluate long context accuracy on 16 tasks in LongBench (Bai et al., 2023). Unless specification, we refer to baseline as a model with non-compressed KV-Cache. See Appendix H for further details on the dataset and settings.
Compression Settings.
We implemented Palu based on the Huggingface library (Wolf et al., 2020). Decomposition of the Key and Value projection layers was performed using the truncation-aware SVD method proposed by SVD-LLM (Wang et al., 2024). Unless otherwise specified, Palu’s results are G-LRD with a group size of 4 (gs-4), with equal rank size for each group. To calculate Fisher information in rank searching, we used 2048 random samples from Wikitext-2, each with a sequence length of 1024. For quantization integration in Palu, we use a simple per-token, asymmetric integer quantization. For evaluation on quantization results, we compare Palu to advanced KV-Cache quantization methods, including Atom (Zhao et al., 2023), KVQaunt (Hooper et al., 2024), and KIVI (Liu et al., 2024b). Refer to Sec. 5 for a brief summary of these methods.
GPU Kernels Implementation.
[…]
4.2 Results with Different Decomposition Granularity
We evaluate perplexity and zero-shot accuracy of Palu with a 50% low-rank compression rate using M-LRD, G-LRD, and J-LRD on Llama-2-7B and Llama-3-8B-Instruct, and present the results in Table 1.
Table 1: Perplexity and zero-shot accuracy of Palu at 50% compression rate.
@@ 4.2 Results with Different Decomposition Granularity
Llama-3-8B-Inst Baseline 8.28 13.01 43.20 75.80 78.62 81.61 56.83 71.90 67.99
J-LRD 9.12 15.90 43.40 73.20 76.50 79.63 51.96 72.45 66.19
G-LRD 10.11 17.87 42.60 70.36 76.06 76.30 48.99 72.38 64.45
M-LRD 12.38 23.02 38.80 63.04 73.67 69.78 42.58 62.51 58.40
@@
As Table 1 shows, for the Llama-2-7B model, Palu’s M-LRD method fails to maintain a low perplexity at a 50% compression rate. In contrast, despite having a high recomputation cost, J-LRD significantly outperforms M-LRD and achieves a 5.62 perplexity on WikiText-2.
For G-LRD, which still maintains a low computation cost, yields a 6.01 perplexity on Wikitext-2, showing a great balance between model accuracy and compression overheads. The same trend is observed in the Llama-3-8B model as well. More results Llama-2-13B can be found in ppendix F.
@@
Similar to the perplexity evaluation, the J-LRD method demonstrates the best performance for the zero-shot accuracy on Llama-2-7B, with only a 0.23% average accuracy degradation. M-LRD method results in the lowest average performance, with a 7.26% drop in accuracy compared to the baseline. In comparison, G-LRD only has a 2.4% average accuracy decline, offering a sweet spot between model accuracy and compression overheads again.
4.3 Results of Quantization Integration
Table 2: Quantization perplexity and KV-Cache size for Llama-2-7B on WikiText-2. For perplexity, sequence length is 4096. KV-Cache size is demonstrated for 128K sequence length.
@@ 4.3 Results of Quantization Integration
Table 2 showcases the impact of quantization on perplexity and KV-Cache size when combined with Palu. With 3-bit quantization, Palu incurs only a slight 0.08 and 0.23 perplexity increase at 30% and 50% low-rank compression rate. These demonstrate a minimal accuracy trade-off for significant compression gains compared to the 16-bit baseline.
Notably, at 2-bit quantization, Palu decisively outperforms the state-of-the-art KVQuant method, reducing perplexity by 1.19 and 0.54, while further slashing memory usage by 30% and 50%. These results establish Palu with quantization as a superior KV-Cache compression method.
4.4 Evaluation on Long Context Datasets
To access Palu’s ability for long-context scenarios, we evaluate baseline, KIVI and Palu’s accuracy on LongBench (Bai et al., 2023). Here we evaluate on Mistral-7B and LongChat-7B models, which have up to 32K context length. We report the average score for each task type separately, as well as the overall average across all 16 tasks. The results are shown in Table 3.
As Table 3 indicates, we find that at a 50% low-rank compression level, Palu is relatively difficult to fully preserve accuracy. However, at a 30% compression level, Palu achieves only a minor average accuracy drop ($<1\%$) compared to the baseline for both models.
Furthermore, Palu can quantize the low-rank latent KV-Cache down to 3 bits, with less than 1% further accuracy degradation. Overall, Palu maintains a strong 40.77% and 34.33% average accuracy for Mistral-7B and LongChat-7B, with an impressive 7.59x compression ratio.
When compared to KIVI (Liu et al., 2024b), Palu achieves a similar accuracy, while having an additional 30% compression rate from low-rank.
Notably, Palu does not require the complex grouped quantization and mixed-precision techniques employed by KIVI, resulting a high inference efficiency (see Sec. 4.5 for details).
4.5 Latency Evaluation
In this section, we provide latency and speedup evaluation, using Llama-2-7B as the base model. We measure latency on a single RTX 4090 GPU and compare Palu to the FP16 and KIVI-4-bit baselines.
We evaluate Palu’s latency at a 50% compression rate, where we set compression rates for key and value to 75% and 25%, respectively. This allocation is based on our observations from the rank allocation results (see Appendix G.3 for details).
For FP16 baseline, we use the default implementation from HuggingFace. For KIVI, we use the CUDA kernels from its official repository. We do not apply FlashAttention (Dao et al., 2022) for an apple-to-apple comparison to our baselines. Due to the small memory capacity of RTX 4090 GPU, we adopt a 4-bit quantization (Frantar et al., 2024) for the weights of all linear layers. Our results are the average of 100 runs.